OpenStudy (anonymous):

EVALUATE ∑5 K=1 K!/K

OpenStudy (anonymous):

\[\begin{align*}\sum_{k=1}^5 \frac{k!}{k}&=\frac{1!}{1}+\frac{2!}{2}+\frac{3!}{3}+\frac{4!}{4}+\frac{5!}{5}\\ &=1+1!+2!+3!+4!\\ &=1+1+2+6+24 \end{align*}\]